# Generative Prompt Internalization

### v2

@@ 1 Introduction
To address this issue, existing methods can be categorized into two approaches. One practical approach only for efficient prompt computation is compressing prompts into external embeddings (Mu et al., 2023; Chevalier et al., 2023; Ge et al., 2024) or compressed text  (Jiang et al., 2023b; Li et al., 2023; Jiang et al., 2024; Pan et al., 2024). However, these methods still require additional tokens, limiting their ability to fully internalize prompts and effectively modify the model’s behavior as intended. Alternatively, fine-tuning (Kim and Rush, 2016; Zou et al., 2024) or distillation (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), have been explored as canonical approaches for internalizing prompts. These methods adjust the language model’s behavior to follow the prompt’s intentions without requiring a prompt during inference.
@@ 2 Related Works
is one of the practical approach to reducing the computational overhead caused by lengthy prompts. In case of the API-based large language models (LLMs) services, systematic caching solutions allow frequently used prompts to be stored between API calls. [^2] https://www.anthropic.com/news/prompt-caching [^3] https://ai.google.dev/gemini-api/docs/caching For users of API-based LLMs, text-based prompt compression methods are proposed, where key segments of long prompts are selected on a token-by-token basis to generate a compressed version of the original prompt (Jiang et al., 2023b; Li et al., 2023; Jiang et al., 2024; Pan et al., 2024). Although this approach reduces prompt length, it often still results in relatively long token sequences, as essential tokens must be retained. On the other hand, embedding-based prompt compression methods generate cached token embeddings for prompts, which can be utilized as a more efficient representation in LLMs (Mu et al., 2023; Chevalier et al., 2023; Ge et al., 2024). These methods offer the advantage of using fewer token embeddings compared to text-based compression but often require modifications to the model architecture, making it challenging to leverage the compressed vectors across different models.
@@ 7.1 Performance
LLMLingua-2 (Pan et al., 2024) exhibits a significant performance drop as the compression rate increases. Although the semantics of the prompt can be inferred, the loss of format crucial for agent tasks leads to substantial performance drops. In the Web Shopping scenario, LLMLingua-2 (Pan et al., 2024) achieves the best performance at a compression rate of 90%. However, compression rates exceeding 30% led to a failure in all tasks. In the case of ICAE (Ge et al., 2024), despite fine-tuning, it struggles to handle the agent application prompts. Many embedding-based compression approaches (Mu et al., 2023; Chevalier et al., 2023), including ICAE (Ge et al., 2024), are primarily optimized for general texts such as articles, and thus exhibit limitations in compressing task-specific information, as required by AgentBench (Liu et al., 2023).
@@ 9 Discussion
Prompt compression approaches, such as LLMLingua-2 (Pan et al., 2024) and ICAE (Ge et al., 2024), effectively compress short instructions without requiring additional training for new prompts (Mu et al., 2023; Chevalier et al., 2023; Ge et al., 2024). However, as shown in Table 1, these methods struggle with long, information-rich application prompts, where key details are often lost in the compression process.
@@ C.3 Criteria of selecting Embedding-based Compression Baseline
Recent embedding-based prompt compression approaches (Mu et al., 2023; Chevalier et al., 2023; Ge et al., 2024) have been developed as a general language models that first cache the given prompt first and then infer user input using compressed prompt embeddings. Gisting (Mu et al., 2023) and AutoCompressor is based on LLaMA-7B (Touvron et al., 2023a) and LLaMA-2-7B (Touvron et al., 2023b), respectively. These models perform significantly poorly on AgentBench (Liu et al., 2023) with scores close to zero. Therefore, we select the baseline ICAE (Ge et al., 2024), which is based on Mistral-7B-Instruct (Jiang et al., 2023a).

### v3

@@ 1 Introduction
To address this issue, existing methods can be categorized into two approaches. One practical approach only for efficient prompt computation is compressing prompts into external embeddings (Chevalier et al., 2023; Mu et al., 2024; Ge et al., 2024) or compressed text  (Jiang et al., 2023b; Li et al., 2023; Jiang et al., 2024; Pan et al., 2024). However, these methods still require additional tokens, limiting their ability to fully internalize prompts and effectively modify the model’s behavior as intended. Alternatively, fine-tuning (Kim and Rush, 2016; Zou et al., 2024) or distillation (Askell et al., 2021; Snell et al., 2022; Choi et al., 2023; Li et al., 2024), have been explored as canonical approaches for internalizing prompts. These methods adjust the language model’s behavior to follow the prompt’s intentions without requiring a prompt during inference.
@@ 2 Related Works
is one of the practical approach to reducing the computational overhead caused by lengthy prompts. In case of the API-based large language models (LLMs) services, systematic caching solutions allow frequently used prompts to be stored between API calls. [^2] https://www.anthropic.com/news/prompt-caching [^3] https://ai.google.dev/gemini-api/docs/caching For users of API-based LLMs, text-based prompt compression methods are proposed, where key segments of long prompts are selected on a token-by-token basis to generate a compressed version of the original prompt (Jiang et al., 2023b; Li et al., 2023; Jiang et al., 2024; Pan et al., 2024). Although this approach reduces prompt length, it often still results in relatively long token sequences, as essential tokens must be retained. On the other hand, embedding-based prompt compression methods generate cached token embeddings for prompts, which can be utilized as a more efficient representation in LLMs (Chevalier et al., 2023; Mu et al., 2024; Ge et al., 2024). These methods offer the advantage of using fewer token embeddings compared to text-based compression but often require modifications to the model architecture, making it challenging to leverage the compressed vectors across different models.
@@ 7.1 Performance
LLMLingua-2 (Pan et al., 2024) exhibits a significant performance drop as the compression rate increases. Although the semantics of the prompt can be inferred, the loss of format crucial for agent tasks leads to substantial performance drops. In the Web Shopping scenario, LLMLingua-2 (Pan et al., 2024) achieves the best performance at a compression rate of 90%. However, compression rates exceeding 30% led to a failure in all tasks. In the case of ICAE (Ge et al., 2024), despite fine-tuning, it struggles to handle the agent application prompts. Many embedding-based compression approaches (Chevalier et al., 2023; Mu et al., 2024), including ICAE (Ge et al., 2024), are primarily optimized for general texts such as articles, and thus exhibit limitations in compressing task-specific information, as required by AgentBench (Liu et al., 2023).
@@ 9 Discussion
Prompt compression approaches, such as LLMLingua-2 (Pan et al., 2024) and ICAE (Ge et al., 2024), effectively compress short instructions without requiring additional training for new prompts (Chevalier et al., 2023; Mu et al., 2024; Ge et al., 2024). However, as shown in Table 1, these methods struggle with long, information-rich application prompts, where key details are often lost in the compression process.
@@ C.3 Criteria of selecting Embedding-based Compression Baseline
Recent embedding-based prompt compression approaches (Chevalier et al., 2023; Mu et al., 2024; Ge et al., 2024) have been developed as a general language models that first cache the given prompt first and then infer user input using compressed prompt embeddings. Gisting (Mu et al., 2024) and AutoCompressor is based on LLaMA-7B (Touvron et al., 2023a) and LLaMA-2-7B (Touvron et al., 2023b), respectively. These models perform significantly poorly on AgentBench (Liu et al., 2023) with scores close to zero. Therefore, we select the baseline ICAE (Ge et al., 2024), which is based on Mistral-7B-Instruct (Jiang et al., 2023a).